OpenStudy (anonymous):

John earns $200 per project for the first five projects, $250 per project for the next 10 projects, and $350 per project for any additional projects he completes. Identify the piecewise function for this situation, and find the amount John will earn for completing 20 Projects

OpenStudy (anonymous):

f(x) = 200 x , x in [0,5] = 1000 + 250 [ x - 10 ] , x in (5, 15] = 2250 + 350 [ x - 15] , 15 < x